OpenStudy (anonymous):

Find the exact value of the equation for 0

OpenStudy (lolacole12):

To solve the equation, we'll apply the formula: sin x = 2 sin x *cos x (1) We'll substitute (1) in the given equation: sin x - sin 2x=0 sin x - 2 sin x *cos x=0 We'll factorize by sin x: sin x*(1- 2cos x)=0 We'll set the factors as 0: sin x=0 or cos x =1/2 sin x =0 and x belongs to [0,2pi) x=(-1)^k*arcsin0 + k*pi x={0,pi} cos x =1/2 x=+/-arccos (1/2)+2*k*pi x={pi/3, 5pi/3}